### v4

@@ 6 Quasigroup Isomorphism
    Bose [Bos63] previously showed that pseudo-STS graphs with $>67$ vertices are STS graphs. Now given a block-incidence graph from a Steiner $2$-design with bounded block size, we can recover the underlying design in $\mathsf{AC^{0}}$ [Lev23, Proposition 4.7]. This yields the desired bound.
@@
Latin square graphs are one of the four families of strongly regular graphs under Neumaier’s classification [Neu79] (the other families being line graphs of Steiner $2$-designs, conference graphs, and graphs whose eigenvalues satisfy the claw bound). Levet [Lev23] previously established an upper bound of $\exists^{\log^{2}n}\mathsf{AC^{0}}$ for isomorphism testing of conference graphs, which is a stronger upper bound than we obtain for Latin square graphs. In contrast, the best known algorithmic runtime for isomorphism testing of conference graphs is $n^{2\log(n)+O(1)}$ due to Babai [Bab80], whereas isomorphism testing of Latin square graphs is known to admit an $n^{\log(n)+O(1)}$-time solution [Mil78].
@@ 7.1 MGS for Groups in $\mathsf{AC}^{1}(\mathsf{L})$
We iterate on this process starting from $G/\text{Soc}(G)$. Note that, as we have computed $\text{Soc}(G)$ from the previous paragraph, we may write down the cosets for $G/\text{Soc}(G)$ in $\mathsf{AC^{0}}$. Furthermore, given a subgroup $H\leq G/\text{Soc}(G)$, we may write down the elements of $H\text{Soc}(G)$ in $\mathsf{AC^{0}}$. By the above, the minimal normal subgroups of a group are computable in $\mathsf{L}$. As there are at most $\log n$ terms in a chief series, we may compute a chief series for $G$ in $\mathsf{AC}^{1}(\mathsf{L})$, as desired. (Recall that we use this notation to mean an $\mathsf{AC}^{1}$ circuit with oracle gates calling a $\mathsf{L}$ oracle, not function composition such as $\mathsf{AC}^{1}\circ\mathsf{L}$.) ∎
@@
Proof of Theorem 7.1.
By Lem. 7.2, we can compute a chief series for $G$ in $\mathsf{AC}^{1}(\mathsf{L})$. So let $N_{1}\triangleleft\cdots\triangleleft N_{k}$ be a chief series $S$ of $G$. Lucchini and Thakkar [LT24] showed that minimum generating sets of $G/N_{i+1}$ have specific structure depending on whether or not $N_{i+1}/N_{i}$ is Abelian. We proceed inductively down $S$ starting from $N_{k-1}$. As $G/N_{k-1}$ is a finite simple group, and hence at most $2$-generated, we can write all $\binom{n}{2}$ possible generating sets in parallel with a single $\mathsf{AC}^{0}$ circuit and test whether each generates the group with a membership test. This can be done in $\mathsf{L}$.
@@
  • Case 1: Suppose that $N=N_{i}/N_{i-1}$ is Abelian. By [LM94, Theorem 4], we have two cases:
@@
    • Case 1b: If Case 1a does not hold, then we necessarily have that $G/N_{i-1}=\langle g_{1},\cdots,g_{d},x\rangle$ for some any non-identity element $x\in N$.
@@
  • Case 2: Suppose instead that $N=N_{i}/N_{i-1}$ is non-Abelian. Then by [LT24, Corollary 13], the following holds. Let $\eta(G)$ denote the number of factors in a chief series with order $|N|$. Let $u=\max\{d,2\}$ and $t=\min\{u,\lceil\frac{8}{5}+\log_{\lvert N\rvert}\eta_{G}(N)\rceil\}$. Then there exist $n_{1},\ldots,n_{t}\in N_{i-1}$ (possibly $n_{1}=\cdots=n_{t}=1$) such that $G/N_{i-1}=\langle g_{1}n_{1},\cdots,g_{t}n_{t},g_{t+1},\cdots,g_{d}\rangle$.
    By [LT24, Corollary 13], there are at most $\lvert N\rvert^{\lceil\frac{8}{5}+\log_{\lvert N\rvert}\eta_{G}(N)\rceil}$ generating sets of this form. As $\log_{|N|}\eta_{G}(N)\in O(1)$, we may write down these generating sets in parallel with a single $\mathsf{AC}^{0}$ circuit and test whether each generate $G/N_{i-1}$ in $\mathsf{L}$ using Membership.
Descending along the chief series in this fashion, we compute quotients $N_{i}/N_{i-1}$ and compute a generating set for $G/N_{i-1}$. The algorithm terminates when we’ve computed a generating set for $G/N_{0}=G$. Since a chief series has $O(\log n)$ terms, this algorithm requires $O(\log n)$ iterations and each iteration is computable in $\mathsf{L}$. Hence, we have an algorithm for MGS in $\mathsf{AC}^{1}(\mathsf{L})$. ∎
Improving upon the $\mathsf{AC}^{1}(\mathsf{L})$ bound on MGS for groups appears daunting. It is thus natural to inquire as to families of groups where MGS is solvable in complexity classes contained within $\mathsf{AC}^{1}(\mathsf{L})$. To this end, we examine the class of nilpotent groups. Arvind & Torán previously established a polynomial-time algorithm for nilpotent groups [AT06, Theorem 7]. We improve their bound as follows.
@@
For a nilpotent group $G$, we can compute $d(G)$ in $\mathsf{L}\cap\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
Proof.
Let $G$ be our input group. Recall that a finite nilpotent group is the direct product of its Sylow subgroups (which by the Sylow theorems, implies that for a given prime $p$ dividing $|G|$, the Sylow $p$-subgroup of $G$ is unique). We can, in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ (using Cor. 3.2), decide if $G$ is nilpotent; and if so, compute its Sylow subgroups. So we write $G=P_{1}\times P_{2}\times\cdots\times P_{\ell}$, where each $P_{i}$ is the Sylow subgroup of $G$ corresponding to the prime $p_{i}$. Arvind & Torán (see the proof of [AT06, Theorem 7]) established that $d(G)=\max_{1\leq i\leq\ell}d(P_{i})$. Thus, it suffices to compute $d(P_{i})$ for each $i\in[\ell]$.
@@
$\mathsf{L}\cap\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ (the latter using Cor. 3.2). We now turn to computing $[P,P]$. Using a membership test, we can compute $[P,P]$ in $\mathsf{L}$. By [Ser97, I.§4 Exercise 5], every element in $[P,P]$ is the product of at most $\log|P|$ commutators. Therefore, we can also decide membership in $[P,P]$ in $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, and so we can write down the elements of $[P,P]$ in $\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
Thus, we may compute $\Phi(P)$ in $\mathsf{L}\cap\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. Given $\Phi(P)$, we may compute $|P/\Phi(P)|$ in $\mathsf{AC^{0}}$. Thus, we may recover $d(P)$ from $|P/\Phi(P)|$ in $\mathsf{AC^{0}}$, by iterated multiplication of the prime divisor $p$ of $|P|$. As the length of the encoding of $p$ is at most $\log|P|$ and we are multiplying $p$ by itself $\log|P|$ times, iterated multiplication is $\mathsf{AC}^{0}$-computable. Thus, in total, we may compute $d(P)$ in $\mathsf{L}\cap\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. It follows that for an arbitrary nilpotent group $G$, we may compute $d(G)$ in $\mathsf{L}\cap\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. ∎
@@
While Prop. 7.3 allows us to compute $d(G)$ for a nilpotent group $G$, the algorithm is non-constructive. It is not clear how to find such a generating set in $\mathsf{L}$. We can, however, provide such a generating set in $\mathsf{AC}^{1}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. Note that this bound is incomparable to $\mathsf{AC}^{1}(\mathsf{L})$. We outline the algorithm here.
The Burnside Basis Theorem provides that for a nilpotent group $G$, (i) every generating set of $G$ projects to a generating set of $G/\Phi(G)$, and (ii) for every generating set $S$ of $G/\Phi(G)$, every lift of $S$ is a generating set of $G$. Furthermore, every minimum generating set of $G$ can be obtained from the Sylow subgroups in the following manner. Write $G=P_{1}\times\cdots\times P_{\ell}$, where $P_{i}$ is the Sylow $p_{i}$-subgroup of $G$. Suppose that $P_{i}=\langle g_{i1},\ldots,g_{ik}\rangle$ (where we may have $g_{ij}=1$ for certain values of $j$). Write $g_{j}=\prod_{i=1}^{\ell}g_{ij}$. As in the proof of [AT06, Theorem 7] we obtain $G=\langle g_{1},\ldots,g_{k}\rangle.$
Given generating sets for $P_{1},\ldots,P_{\ell}$, we may in $\mathsf{L}\cap\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ recover a generating set for $G$. Thus, it suffices to compute a minimum generating set $P/\Phi(P)\cong(\mathbb{Z}/p\mathbb{Z})^{d(P)}$, where $P$ is a $p$-group. Note that we may handle each Sylow subgroup of $G$ in parallel. To compute a minimum generating set of $P/\Phi(P)$, we use the generator enumeration strategy. As $P/\Phi(P)$ is Abelian, we may check in $\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ Cor. 3.6 whether a set of elements generates the group. As $d(P)\leq\log|P|$, we have $\log|P|$ steps where each step is $\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$-computable. Thus, we may compute a minimum generating set for $P$ in $\mathsf{AC}^{1}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$, as desired.
@@ 7.2 MGS for Quasigroups
In the paper in which they introduced (polylog-)limited nondeterminism, Papadimitriou and Yannakakis conjectured that MGS for quasigroups was $\exists^{\log^{2}n}\mathsf{P}$-complete [PY96, after Thm. 7]. While they did not specify the type of reductions used, it may be natural to consider polynomial-time many-one reductions. Theorem 7.5 refutes two versions of their conjecture under other kinds of reductions, that are incomparable to polynomial-time many-one reductions: $\mathsf{quasiAC^{0}}$ reductions unconditionally and polylog-space reductions conditionally. We note that their other $\exists^{\log^{2}n}\mathsf{P}$-completeness result in the same section produces a reduction that in fact can be done in logspace and (with a suitable, but natural, encoding of the gates in a circuit) also in $\mathsf{AC^{0}}$, so our result rules out any such reduction for MGS. (We also note: assuming $\mathsf{EXP}\neq\mathsf{PSPACE}$, showing that this problem is complete under polynomial-time reductions would give a separation between poly-time and log-space reductions, an open problem akin to $\mathsf{P}\neq\mathsf{L}$.)
@@
Proof.
(a) Thm. 7.5 (a) for MGS, resp. Thm. 6.1 for Quasigroup Isomorphism, place both problems into classes that are contained in $\mathsf{quasiAC^{0}}$ by Fact 2.4. If either problem were $\exists^{\log^{2}n}\mathsf{P}$-complete under quasi-polynomial size, $o(\log n/\log\log n)$-depth, $\mathsf{AC}[p]$ circuits, then, since $\textsc{Parity}\in\mathsf{P}\subseteq\exists^{\log^{2}n}\mathsf{P}$, we would get such circuits for Parity, which do not exist [Raz93, Smo87] (Smolensky’s argument yields a minimum size of $\exp(\Omega(n^{1/(2d)}))$ for depth-$d$ circuits, which is super-polynomial when $d\in o(\log n/\log\log n)$).
(b) Both MGS for quasigroups and Quasigroup Isomorphism are in $\mathsf{DSPACE}(\log^{2}n)$ by Thm. 7.5, resp. [CTW13]. The closure of $\mathsf{DSPACE}(\log^{2}n)$ under poly-log space reductions is contained in $\mathsf{polyL}=\bigcup_{k\geq 0}\mathsf{DSPACE}(\log^{k}n)$. If either of these two quasigroup problems were complete for $\exists^{\log^{2}n}\mathsf{P}$ under polylog-space Turing reductions, we would get $\exists^{\log^{2}n}\mathsf{P}\subseteq\mathsf{polyL}$. Under the latter assumption, by a straightforward padding argument, we now show that $\mathsf{EXP}=\mathsf{PSPACE}$.
Let $L\in\mathsf{EXP}$; let $k$ be such that $L\in\mathsf{DTIME}(2^{n^{k}}+k)$. Define $L_{pad}=\{(x,1^{2^{|x|^{k}}+k}):x\in L\}$. By construction, $L_{pad}\in\mathsf{P}$. Let us use $N$ to denote the size of the input to $L_{pad}$, that is, $N=2^{n^{k}}+k+n$. By assumption, we thus have $L_{pad}\in\mathsf{polyL}$. Suppose $\ell$ is such that $L_{pad}\in\mathsf{DSPACE}(\log^{\ell}N)$. We now give a $\mathsf{PSPACE}$ algorithm for $L$. In order to stay within polynomial space, we cannot write out the padding $1^{2^{n^{k}}+k}$ explicitly. What we do instead is simulate the $\mathsf{DSPACE}(\log^{\ell}N)$ algorithm for $L_{pad}$ as follows. Whenever the head on the input tape would move off the $x$ and into the padding, we keep track of its index into the padding, and the simulation responds as though the tape head were reading a 1. When the tape head moves right the index increases by 1, when it moves left it decreases by 1, and if the index is zero and the tape head moves left, then we move the tape head onto the right end of the string $x$. The index itself is a number between 00 and $2^{n^{k}}+k$, so can be stored using only $O(n^{k})$ bits. The remainder of the $L_{pad}$ algorithm uses only $O(\log^{\ell}N)=O(n^{k\ell})$ additional space, thus this entire algorithm uses only a polynomial amount of space, so $L\in\mathsf{PSPACE}$, and thus $\mathsf{EXP}=\mathsf{PSPACE}$. ∎
Now we return to establishing the main result of this section, Thm. 7.5. To establish Thm. 7.5 (a) and (b), we will crucially leverage the Membership for quasigroups problem. To this end, we will first establish the following.
@@
Thm. 7.7 immediately yields the following corollary.
@@
The proofs of Theorem 7.7 and Theorem 7.5 rely crucially on the following adaption of the Babai–Szemerédi Reachability Lemma [BS84, Theorem 3.1] to quasigroups. We first generalize the notion of a straight-line program for groups [BS84] to SLPs for quasigroups. We follow the same strategy as in the proof of [BS84, Theorem 3.1], but there are some subtle, yet crucial, modifications due to the fact that quasigroups are non-associative and need not posses an identity element.
Let $X$ be a set of generators for a quasigroup $G$. We call a sequence of elements $g_{1},\ldots,g_{\ell}\in G$ a straight-line program (SLP for short) if each $g_{i}$ ($i\in[\ell]$) either belongs to $X$, or is of the form or $g_{j}g_{k}$, $g_{j}\backslash g_{k}$, or $g_{j}/g_{k}$ for some $j,k<i$ (where $g_{j}\backslash g_{k}$, resp. $g_{j}/g_{k}$, denotes the quasigroup division as defined in Section 2.1). An SLP is said to compute or generate a set $S$ (or an element $g$) if $S\subseteq\{g_{1},\dots,g_{\ell}\}$ (resp. $g\in\{g_{1},\dots,g_{\ell}\}$).
@@
  • Otherwise, $L(i)$ is not a sub-quasigroup; hence, there exist $g,h\in L(i)$ with one of $gh$, $g/h$, or $g\backslash h\not\in L(i)$. We choose $z_{i+1}$ to be one of these $gh$, $g/h$, or $g\backslash h\not\in L(i)$.
Next, we first claim that $|K(i+1)|=2\cdot|K(i)|$. Note that $K(i+1)=K(i)\cup K(i)z_{i+1}$ by definition. As right-multiplication by a fixed element is a bijection in a quasi-group, it suffices to show that $K(i)\cap K(i)z_{i+1}=\emptyset$. So, suppose that there exists some $a\in K(i)\cap K(i)z_{i+1}$. Then $a=gz_{i+1}$ for some $g\in K(i)$. Hence, $z_{i+1}=g\backslash a$, contradicting $z_{i+1}\not\in L(i)$, since both $a$ and $g$ are in $K(i)$.

### v5

@@ 6 Quasigroup Isomorphism
    Bose [15] previously showed that pseudo-STS graphs with >67>67 vertices are STS graphs. Now given a block-incidence graph from a Steiner 22-design with bounded block size, we can recover the underlying design in $\mathsf{AC^{0}}$ [37, Proposition 4.7]. This yields the desired bound.
@@
Latin square graphs are one of the four families of strongly regular graphs under Neumaier’s classification [44] (the other families being line graphs of Steiner $2$-designs, conference graphs, and graphs whose eigenvalues satisfy the claw bound). Levet [37] previously established an upper bound of $\exists^{\log^{2}n}\mathsf{AC^{0}}$ for isomorphism testing of conference graphs, which is a stronger upper bound than we obtain for Latin square graphs. In contrast, the best known algorithmic runtime for isomorphism testing of conference graphs is $n^{2\log(n)+O(1)}$ due to Babai [6], whereas isomorphism testing of Latin square graphs is known to admit an $n^{\log(n)+O(1)}$-time solution [43].
@@ 7.1 MGS for Groups in $\mathsf{AC}^{1}(\mathsf{L})$
We iterate on this process starting from $G/\text{Soc}(G)$. Note that, as we have computed $\text{Soc}(G)$ from the previous paragraph, we may write down the cosets for $G/\text{Soc}(G)$ in $\mathsf{AC^{0}}$. Furthermore, given a subgroup $H\leq G/\text{Soc}(G)$, we may write down the elements of $H\text{Soc}(G)$ in $\mathsf{AC^{0}}$. By the above, the minimal normal subgroups of a group are computable in $\mathsf{L}$. As there are at most $\log n$ terms in a chief series, we may compute a chief series for $G$ in $\mathsf{AC}^{1}(\mathsf{L})$, as desired. (Recall that we use this notation to mean an $\mathsf{AC}^{1}$ circuit with oracle gates calling a $\mathsf{L}$ oracle, not function composition such as $\mathsf{AC}^{1}\circ\mathsf{L}$.)
@@
Proof 7.4 (Proof of Theorem 7.1).
By Lem. 7.2, we can compute a chief series for $G$ in $\mathsf{AC}^{1}(\mathsf{L})$. So let $N_{1}\triangleleft\cdots\triangleleft N_{k}$ be a chief series $S$ of $G$. Lucchini and Thakkar [40] showed that minimum generating sets of $G/N_{i+1}$ have specific structure depending on whether or not $N_{i+1}/N_{i}$ is Abelian. We proceed inductively down $S$ starting from $N_{k-1}$. As $G/N_{k-1}$ is a finite simple group, and hence at most $2$-generated, we can write all $\binom{n}{2}$ possible generating sets in parallel with a single $\mathsf{AC}^{0}$ circuit and test whether each generates the group with a membership test. This can be done in $\mathsf{L}$.
@@
  • Case 1: Suppose that $N=N_{i}/N_{i-1}$ is Abelian. By [39, Theorem 4], we have two cases:
@@
    • Case 1b: If Case 1a does not hold, then we necessarily have that
      $G/N_{i-1}=\langle g_{1},\cdots,g_{d},x\rangle$ for any non-identity element $x\in N$.
@@
  • Case 2: Suppose instead that $N=N_{i}/N_{i-1}$ is non-Abelian. Then by [40, Corollary 13], the following holds. Let $\eta_{G}(N)$ denote the number of factors in a chief series with order $|N|$. Let $u=\max\{d,2\}$ and $t=\min\{u,\lceil\frac{8}{5}+\log_{\lvert N\rvert}\eta_{G}(N)\rceil\}$. Then there exist $n_{1},\ldots,n_{t}\in N_{i-1}$ (possibly $n_{1}=\cdots=n_{t}=1$) such that $G/N_{i-1}=\langle g_{1}n_{1},\cdots,g_{t}n_{t},g_{t+1},\cdots,g_{d}\rangle$.
    By [40, Corollary 13], there are at most $\lvert N\rvert^{\lceil\frac{8}{5}+\log_{\lvert N\rvert}\eta_{G}(N)\rceil}$ generating sets of this form. As $\log_{|N|}\eta_{G}(N)\in O(1)$, we may write down these generating sets in parallel with a single $\mathsf{AC}^{0}$ circuit and test whether each generate $G/N_{i-1}$ in $\mathsf{L}$ using Membership.
Descending along the chief series in this fashion, we compute quotients $N_{i}/N_{i-1}$ and compute a generating set for $G/N_{i-1}$. The algorithm terminates when we’ve computed a generating set for $G/N_{0}=G$. Since a chief series has $O(\log n)$ terms, this algorithm requires $O(\log n)$ iterations and each iteration is computable in $\mathsf{L}$. Hence, we have an algorithm for MGS in $\mathsf{AC}^{1}(\mathsf{L})$.
Improving upon the $\mathsf{AC}^{1}(\mathsf{L})$ bound on MGS for groups appears daunting. It is thus natural to inquire as to families of groups where MGS is solvable in complexity classes contained within $\mathsf{AC}^{1}(\mathsf{L})$. To this end, we examine the class of nilpotent groups. Arvind & Torán previously established a polynomial-time algorithm for nilpotent groups [3, Theorem 7]. We improve their bound as follows.
@@
For a nilpotent group $G$, we can compute $d(G)$ in
$\mathsf{L}\cap\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))).$
Proof 7.6.
Let $G$ be our input group. Recall that a finite nilpotent group is the direct product of its Sylow subgroups (which by the Sylow theorems, implies that for a given prime $p$ dividing $|G|$, the Sylow $p$-subgroup of $G$ is unique). We can, in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ (using Cor. 3.3), decide if $G$ is nilpotent; and if so, compute its Sylow subgroups. So we write $G=P_{1}\times P_{2}\times\cdots\times P_{\ell}$, where each $P_{i}$ is the Sylow subgroup of $G$ corresponding to the prime $p_{i}$. Arvind & Torán (see the proof of [3, Theorem 7]) established that $d(G)=\max_{1\leq i\leq\ell}d(P_{i})$. Thus, it suffices to compute $d(P_{i})$ for each $i\in[\ell]$.
@@
$\mathsf{L}\cap\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ (the latter using Cor. 3.3). We now turn to computing $[P,P]$. Using a membership test, we can compute $[P,P]$ in $\mathsf{L}$. By [51, I.§4 Exercise 5], every element in $[P,P]$ is the product of at most $\log|P|$ commutators. Therefore, we can also decide membership in $[P,P]$ in $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, and so we can write down the elements of $[P,P]$ in $\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
Thus, we may compute $\Phi(P)$ in $\mathsf{L}\cap\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. Given $\Phi(P)$, we may compute $|P/\Phi(P)|$ in $\mathsf{AC^{0}}$. Thus, we may recover $d(P)$ from $|P/\Phi(P)|$ in $\mathsf{AC^{0}}$, by iterated multiplication of the prime divisor $p$ of $|P|$. As the length of the encoding of $p$ is at most $\log|P|$ and we are multiplying $p$ by itself $\log|P|$ times, iterated multiplication is $\mathsf{AC}^{0}$-computable. Thus, in total, we may compute $d(P)$ in $\mathsf{L}\cap\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. It follows that for an arbitrary nilpotent group $G$, we may compute $d(G)$ in $\mathsf{L}\cap\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
@@
While Prop. 7.5 allows us to compute $d(G)$ for a nilpotent group $G$, the algorithm is non-constructive. It is not clear how to find such a generating set in $\mathsf{L}$. We can, however, compute such a generating set in $\mathsf{AC}^{1}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. Note that this bound is incomparable to $\mathsf{AC}^{1}(\mathsf{L})$. We outline the algorithm here.
The Burnside Basis Theorem provides that for a nilpotent group $G$, (i) every generating set of $G$ projects to a generating set of $G/\Phi(G)$, and (ii) for every generating set $S$ of $G/\Phi(G)$, every lift of $S$ is a generating set of $G$. Furthermore, every minimum generating set of $G$ can be obtained from the Sylow subgroups in the following manner. Write $G=P_{1}\times\cdots\times P_{\ell}$, where $P_{i}$ is the Sylow $p_{i}$-subgroup of $G$. Suppose that $P_{i}=\langle g_{i1},\ldots,g_{ik}\rangle$ (where we may have $g_{ij}=1$ for certain values of $j$). Write $g_{j}=\prod_{i=1}^{\ell}g_{ij}$. As in the proof of [3, Theorem 7] we obtain $G=\langle g_{1},\ldots,g_{k}\rangle.$
Given generating sets for $P_{1},\ldots,P_{\ell}$, we may in $\mathsf{L}\cap\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ recover a generating set for $G$. Thus, it suffices to compute a minimum generating set for $P/\Phi(P)\cong(\mathbb{Z}/p\mathbb{Z})^{d(P)}$, where $P$ is a $p$-group. Note that we may handle each Sylow subgroup of $G$ in parallel. To compute a minimum generating set of $P/\Phi(P)$, we use the generator enumeration strategy. As $P/\Phi(P)$ is Abelian, we may check in $\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ (by Cor. 3.10) whether a set of elements generates the group. As $d(P)\leq\log|P|$, we have $\log|P|$ steps where each step is $\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$-computable. Thus, we may compute a minimum generating set for $P$ in $\mathsf{AC}^{1}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$, as desired.
@@ 7.2 MGS for Quasigroups
In the paper in which they introduced (polylog-)limited nondeterminism, Papadimitriou and Yannakakis conjectured that MGS for quasigroups was $\exists^{\log^{2}n}\mathsf{P}$-complete [45, after Thm. 7]. While they did not specify the type of reductions used, it may be natural to consider polynomial-time many-one reductions. Theorem 7.8 refutes two versions of their conjecture under other kinds of reductions, that are incomparable to polynomial-time many-one reductions: $\mathsf{quasiAC^{0}}$ reductions unconditionally and polylog-space reductions conditionally. We note that their other $\exists^{\log^{2}n}\mathsf{P}$-completeness result in the same section produces a reduction that in fact can be done in logspace and (with a suitable, but natural, encoding of the gates in a circuit) also in $\mathsf{AC^{0}}$, so our result rules out any such reduction for MGS. (We also note: assuming $\mathsf{EXP}\neq\mathsf{PSPACE}$, showing that this problem is complete under polynomial-time reductions would give a separation between poly-time and log-space reductions, an open problem akin to $\mathsf{P}\neq\mathsf{L}$.)
@@
Proof 7.10.
(a) Thm. 7.8 (a) for MGS, resp. Thm. 6.1 for Quasigroup Isomorphism, place both problems into classes that are contained in $\mathsf{quasiAC^{0}}$ by Fact 1. If either problem were $\exists^{\log^{2}n}\mathsf{P}$-complete under quasi-polynomial size, $o(\log n/\log\log n)$-depth, $\mathsf{AC}[p]$ circuits, then, since $\textsc{Parity}\in\mathsf{P}\subseteq\exists^{\log^{2}n}\mathsf{P}$, we would get such circuits for Parity, which do not exist [46, 52] (Smolensky’s argument yields a minimum size of $\exp(\Omega(n^{1/(2d)}))$ for depth-$d$ circuits, which is super-polynomial when $d\in o(\log n/\log\log n)$).
(b) Both MGS for quasigroups and Quasigroup Isomorphism are in $\mathsf{DSPACE}(\log^{2}n)$ by Thm. 7.8, resp. [17]. The closure of $\mathsf{DSPACE}(\log^{2}n)$ under poly-log space reductions is contained in $\mathsf{polyL}=\bigcup_{k\geq 0}\mathsf{DSPACE}(\log^{k}n)$. If either of these two quasigroup problems were complete for $\exists^{\log^{2}n}\mathsf{P}$ under polylog-space Turing reductions, we would get $\exists^{\log^{2}n}\mathsf{P}\subseteq\mathsf{polyL}$. Under the latter assumption, by a straightforward padding argument, we now show that $\mathsf{EXP}=\mathsf{PSPACE}$.
Let $L\in\mathsf{EXP}$; let $k$ be such that $L\in\mathsf{DTIME}(2^{n^{k}}+k)$. Define $L_{pad}=\{(x,1^{2^{|x|^{k}}+k}):x\in L\}$. By construction, $L_{pad}\in\mathsf{P}$. Let us use $N$ to denote the size of the input to $L_{pad}$, that is, $N=2^{n^{k}}+k+n$. By assumption, we thus have $L_{pad}\in\mathsf{polyL}$. Suppose $\ell$ is such that $L_{pad}\in\mathsf{DSPACE}(\log^{\ell}N)$. We now give a $\mathsf{PSPACE}$ algorithm for $L$. In order to stay within polynomial space, we cannot write out the padding $1^{2^{n^{k}}+k}$ explicitly. What we do instead is simulate the $\mathsf{DSPACE}(\log^{\ell}N)$ algorithm for $L_{pad}$ as follows. Whenever the head on the input tape would move off the $x$ and into the padding, we keep track of its index into the padding, and the simulation responds as though the tape head were reading a 1. When the tape head moves right the index increases by 1, when it moves left it decreases by 1, and if the index is zero and the tape head moves left, then we move the tape head onto the right end of the string $x$. The index itself is a number between 0 and $2^{n^{k}}+k$, so can be stored using only $O(n^{k})$ bits. The remainder of the $L_{pad}$ algorithm uses only $O(\log^{\ell}N)=O(n^{k\ell})$ additional space, thus this entire algorithm uses only a polynomial amount of space, so $L\in\mathsf{PSPACE}$, and thus $\mathsf{EXP}=\mathsf{PSPACE}$.
Now we return to establishing the main result of this section, Thm. 7.8. To establish Thm. 7.8 (a) and (b), we will crucially leverage the Membership for quasigroups problem. To this end, we will first establish the following.
@@
Thm. 7.11 immediately yields the following corollary.
@@
The proofs of Theorem 7.11 and Theorem 7.8 rely crucially on the following adaption of the Babai–Szemerédi Reachability Lemma [7, Theorem 3.1] to quasigroups. We first generalize the notion of a straight-line program for groups [7] to SLPs for quasigroups. We follow the same strategy as in the proof of [7, Theorem 3.1], but there are some subtle, yet crucial, modifications due to the fact that quasigroups are non-associative and need not possess an identity element.
[…]
@@
[…]